Generate the code to convert the following data frame to wide format. grpsexmeanL sdLmeanR SdR 1A F 0.220.110.340.08 2 A MI 0.47033 0.57033 B F 0.330.11 0.400.0 4B M 0.550.30.650.27 The result should look like the following display. grpF.meanLF.meanRF.sdLF.sdRM.meanLM.meanRM.sdLM.sdR 0.57 0.33 0.33 0.65 0.31 0.27 1A 022 0.34 0.110.08 0.47 0. (), :多 0.40 0.11 0.0 / Hint: use gather() in conjunction with spread() and unite()
Answer #1

#Loading Libraries
library(tidyr)

#Creating the dataframe
data <- data.frame(grp = c("A", "A", "B", "B"),
sex = c("F", "M", "F", "M"),
meanL = c(0.22, 0.47, 0.33, 0.55),
sdL = c(0.11, 0.33, 0.11, 0.31),
meanR = c(0.34, 0.57, 0.4, 0.65),
sdR = c(0.08, 0.33, 0.07, 0.27))


#Producing the result
new_data <- data %>%
gather("statistics", "value", -c("grp", "sex")) %>%
unite(sex.statistics, c("sex", "statistics"), sep = ".") %>%
arrange(sex.statistics) %>%
spread(key = "sex.statistics", value = "value")

**If result does not match please comment